OpenStudy (zaxoanl):

find the volume of the solid generated by revolving the region bounded by the graphs of the equations about the x-axis

OpenStudy (zaxoanl):

\[y=1/x \] \[x \ge1\] \[y=0\]

OpenStudy (zaxoanl):

please how me why the area of the surface is infinite

OpenStudy (tkhunny):

The integrals look like what? Let's see the Volume Integral...

OpenStudy (zaxoanl):

i used the shell method \[2\pi \int\limits_{1}^{\infty} (1+x)(1-1/x)dx\]

OpenStudy (zaxoanl):

the answer came out to be pi not sure if thats correct or not and i don't how to show the surface area is infinite

zepdrix (zepdrix):

Hmmmm surface area isn't working out nicely for me :( I'll show you what I've got so far...

zepdrix (zepdrix):